OpenStudy (anonymous):

Why does this make no sense to me at all If you have Sqrt of 1-(.50)^2 the calcultor says its .87 How is that possible The sqrt of 1 is 1 1-.5 is .5 Can someone explain this

OpenStudy (anonymous):

\[\sqrt{1-(.50)^2}\]

OpenStudy (anonymous):

be careful when entering it

OpenStudy (lalaly):

use brackets

OpenStudy (anonymous):

i did and the answer in my math book is .87

OpenStudy (anonymous):

if type in sqrt (1-(.50)^2) just liek that it says .87 how does that make sense at all The sqrt of 1 is 1 and 1 -,5 is not .87

OpenStudy (lalaly):

first do 1-.5^2 then take the sqrt of the answer i tried it

OpenStudy (anonymous):

like lalaly said, you have to do the stuff under the sqrt first...

OpenStudy (lalaly):

In addition and subtraction \[\sqrt{1-0.5^2} \cancel{=} \sqrt{1}-\sqrt{0.5^2}\]

OpenStudy (anonymous):

the sqrt of .5 is not .87 its .7 so what is going on

OpenStudy (lalaly):

as i said u have to find whats under the root then find the square root, u cant take the sqrt of each number individually

OpenStudy (anonymous):

nevermind im a moron ugh i hate having breaks at school i come back from break and forget everything i KNOW how to do sorry thanks

OpenStudy (lalaly):

its ok... goodluck

OpenStudy (anonymous):

You're not a moron. Everyone forgets stuff. That's why we're here to help. :)

OpenStudy (anonymous):

thanks guys p.s i hate trig lol